A steel ball rolls with constant velocity across a tabletop .950m high. It rolls off and hits the ground +.352m horizontally from the edge of the table. How fast was the ball rolling?
1
Expert's answer
2018-09-06T16:41:08-0400
The horizontal distance
L=vt
The vertical distance
H=(gt^2)/2
Thus
t=√(2H/g)
Finally
v=L/t=L/√(2H/g)=0.352/√((2×0.952)/9.81)=0.8 m/s